OpenStudy (anonymous):

A test has 5 questions. For each question the answer is either correct or incorrect. How many outcomes have more than one correct answer?

OpenStudy (anonymous):

Do you know the right answer?

OpenStudy (anonymous):

Well I know that the possible outcomes for the test is 32

OpenStudy (anonymous):

And if I know that if at least one of the answers is correct it's 31/32

OpenStudy (anonymous):

"How many outcomes have more than one correct answer?" -- This is alluding towards combinatorics not probability.

OpenStudy (anonymous):

Yes it is

OpenStudy (anonymous):

I thought it was 32-6 equal to or less than one correct answer but I could be wrong

OpenStudy (anonymous):

@987poi: Do you know the answer for this question?

OpenStudy (anonymous):

No I don't

OpenStudy (anonymous):

I think it should be \( 2^5 -(5+1) = 16 \), but I will wait for @Zarkon or someone else to corroborate.

OpenStudy (anonymous):

Typo: Read \( 2^5 -(5+1) = 16 \) as \( 2^5 -(5+1) = 26 \)

OpenStudy (anonymous):

Thank you for your input

OpenStudy (anonymous):

Glad to help :)

OpenStudy (zarkon):

26 is correct

OpenStudy (anonymous):

Thanks @Zarkon